OpenStudy (anonymous):

At 1 atm, how much energy is required to heat 49.0 g of H2O(s) at -20 degrees C to H2O(g) at 149.0 degrees C?

OpenStudy (anonymous):

First write down everything you will need. The specific heat capacity for ice, water, steam and the heats of melting and vaporizaton.\[C_s\ water = 4.184j/g K \]\[C_s\ ice = 1.99j/g K \]\[C_s\ steam = 2.11j/g K \]\[\Delta H melting = 334j/g K \]\[ \Delta H = 2260j/g \] Then we use our handy dandy equation: \[q=mass(g)\times C_s \times \Delta T\] and plug it in for each one. Add them all up and you're done! Try that and i'll be back on in a minute to answer any questions.